OpenStudy (anonymous):

How do I do this problem... Which value of b would make 16x^2-bx+25 a perfect square trinomial?

OpenStudy (anonymous):

must be \((4x+5)^2=16x^2+40x+25\) so \(b=40\)

OpenStudy (anonymous):

ooops i ignored the minus sign must be \(16x^2-40x+25\) so \(b=-40\)

OpenStudy (anonymous):

Okay thank you! I think I get the concept!

OpenStudy (anonymous):

yw